John was studying how far the UMN students travelled to take the class on campus. Let W be the distance travelled (measured in miles). A sample of size 25 was collected. He calculated the following statistics: John wanted to run a hypothesis test with the null and alternative hypothesis as below: Ho : μ = 5 Note that μ is the mean distance traveled by UMN students to come to campus. Please calculate the test statistic for the hypothesis test. Please round the answer to the nearest hundredth.
Answer #1

The given information is as shown below: Sample size, n= 25 W = 3.8 25 E(w, -w) = 23 i=1 The null and alternative hypotheses

Thus, 3.8 – 5 0.9789/25 -1.2 0.19578 =-6.12933 x -6.129 is -6.129 Therefore, the t-sta The degrees of freedom is, df =n-1 = 2